# Rydberg Atomic Receiver: Next Frontier of Wireless Communications

### v2

@@ Abstract
Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic (EM) wave measurement by harnessing the electron transition phenomenon of Rydberg atoms. Operating at the quantum scale, such receivers have the potential to breakthrough the performance limit of classic receivers, sparking a revolution in physical-layer wireless communications. The objective of this paper is to offer insights into RARE-empowered communications. We first provide a comprehensive introduction to the fundamental principles of RAREs. Then, a thorough comparison between RAREs and classic receivers is conducted in terms of the antenna size, sensitivity, bandwidth, and hardware cost. Subsequently, we overview the recent progresses in RARE-aided wireless communications, covering the frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques. Moreover, the unique application of RARE in multiband sensing and communication is introduced. Finally, we conclude by providing promising research directions.
@@ I Introduction
In current communication systems, receivers detect wireless signals by using an antenna to transform incident EM waves into electric currents [1]. This mechanism, however, presents insurmountable limitations to physical-layer (PHY) wireless communications. Firstly, the front-end circuits of a receiver, like amplifiers and mixers, introduce thermal noise into data currents, restricting the receiver’s sensitivity limit. In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its working bandwidth to a narrow range around the carrier frequency. To detect signals from sub-6G to THz, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band. Consequently, these challenges constrain the ultimate performance of wireless communications [2].
@@
Recently, at the intersection of quantum sensing and wireless communication, Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3]. Rydberg atoms are highly-excited atoms with outer electrons far from the nucleus. Due to the large electric dipole moments, Rydberg atoms can strongly interact with incident EM waves, triggering electron transitions between energy levels. As a result, RAREs can accurately detect wireless signals by monitoring changes on the atoms’ quantum states resulting from electron transitions [2]. Representing a paradigm shift in receiver architecture to embrace quantum sensing, RAREs are gaining increasing attentions due to their quantum advantages. For instance, a RARE is immune to thermal noise as it involves no front-end circuits. Hence, its sensitivity limit can far exceed those of classic counterparts [4]. Moreover, a single RARE is able to interact with a large number of frequencies ranging from Megahertz (MHz) to THz band because of the abundant energy levels in Rydberg atoms [5]. These quantum advantages make RAREs a promising technology to revolutionize PHY communications in diverse applications, such as weak signal detection, coverage enhancement, and full-frequency sensing.
@@
The objective of this paper is to offer a beginner’s guide to the next-generation wireless communications aided by RAREs. To this end, we first comprehensively introduce the fundamental features of Rydberg atoms and the key modules of a RARE. Then, the performance of RAREs is compared with classic receivers in antenna size, sensitivity, bandwidth, and hardware cost. Subsequently, we overview the state-of-the-art RARE techniques for wireless communications, covering frequency-division multiplexing (FDM), multiple-input-multiple-output (MIMO), wireless sensing, and quantum many-body techniques. Then, the potential of RARE in multiband sensing and communication services is revealed using experimental results. Last, we highlight promising future research directions on RARE from the PHY communication perspective.
@@ II-A3 Properties of Rydberg atom
Given the high principal quantum number $n$, Rydberg atoms exhibit useful properties for sensing EM waves. First, the radius of an Rydberg atom is large due to its dependence on $n^{2}$ [2]. For instance, the atomic radius grows from $150a_{0}$ to $10^{4}a_{0}$ as $n$ increases from $10$ to $100$, with $a_{0}=5.29\times 10^{-11}{\rm m}$ being the Bohr radius. The substantial increase in radius endows the Rydberg atom with a large electric dipole moment, thus making it very sensitive to external EM wave. Second, the transition frequency between adjacent Rydberg states is abundant as the energy level spacing scales as $n^{-3}$ [3]. For low energy levels ($n<20$), the transition frequency can reach the visible light frequency ($>100$ THz). While for Rydberg states ($n>20$), the transition frequencies fall within the range of MHz to THz [4], exactly covering all the typical frequency bands used in PHY communications.
@@ II-B2 EIT based quantum measurement
Figure 2: EIT for measuring AM/FM/PM signals. For AM and PM detection, the detuning $\delta$ is zero, while for FM detection, the detuning $\delta$ is nonzero.
@@
  • The $(1,2)$-th element of the final quantum state, $\rho_{12}$, determines the rate at which Rydberg atoms absorb photons from the probe laser. The absorption loss of the probe-laser scales exponentially with $\rho_{12}$ [6]. Thereby, it is feasible to deduce the value of $\rho_{12}$ by measuring the output power of the probe laser using a photodetector (PD), as depicted in Fig. 1.
@@ II-B3 Receive signal processing
So far, we have discussed the mechanism of detecting AM/FM/PM wireless signals using RAREs. Recent studies have also reported the detection of multi-frequency signals [10], multi-user signals [8], and target vibrations [4], all based on the principles we have covered and will be elaborated in Section IV.
@@
The RARE technique represents a paradigm shift in wireless signal detection from using electronic circuits to embracing quantum sensing. This gives rise to a range of unique features for RARE-enabled PHY communications, which are introduced in this section.
@@ III-A Antenna Size
According to the Chu limit, a lossless dipole antenna in a classic receiver has a size of half-wavelength, $\frac{\lambda}{2}$. This wavelength-dependent size leads to bulky receivers at low frequencies, e.g., exceeding 10 cm when $f<1$ GHz. In contrast, the vapor cells (i.e., antennas of RAREs) have wavelength-independent sizes, e.g., typically fixed as $1$ cm in diameter [6]. Consequently, RAREs can have electrically small antennas, offering significant advantages especially for low-frequency applications.
@@ III-C Bandwidth
In existing RARE studies, the term ”bandwidth” is often used to represent two distinct metrics. Firstly, it is utilized to define the range of frequencies that can trigger electron transitions of Rydberg atoms. This concept is commonly referred to as the “detectable frequency range” of a RARE in the literature. Secondly, it is also associated with the “instantaneous bandwidth”, which denotes the maximum data capacity that the receiver device can measure at any given moment. This term aligns with the standard definition of ”bandwidth” in the field of communication engineering. Typically, a RARE is characterized by a broad detectable frequency range but a narrow instantaneous bandwidth.
@@ III-C2 Instantaneous bandwidth
Restricted by the long response time for an EIT process to become steady, the instantaneous bandwidth of a RARE is typically narrower than 10 MHz [7]. This bandwidth is comparable to that of 4G LTE systems (1$\sim$20 MHz). Recent efforts are endeavored to increase the RARE’s bandwidth. For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time (see [139] of [2]). This technique expands the RARE’s bandwidth to 100 MHz, approaching the bandwidth of sub-6G systems specified by 5G standards (i.e., 20$\sim$100 MHz). To date, continuously improving RARE’s instantaneous bandwidth remains a hot topic.
@@ IV State-of-the-art RARE Techniques
Recently, there have been studies to develop advanced RARE techniques to enhance the communication and sensing performance. This section provides an overview of the state-of-the-art designs.
@@ IV-A Frequency-division Multiplexed RAREs
The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference. The advancement of FDM-RARE is thus regarded as the key enabler of RARE-empowered wideband communications. The state-of-the-art FDM-RAREs can be separated into two categories: standard and multi-band FDM-RAREs.
@@ IV-A1 Standard FDM-RARE
IV-A2 Multiband FDM-RARE
The multiband FDM-RARE leverages the broad detectable frequency range of a RAREs to serve as an integrated full-frequency communication platform. On this platform, signal frequencies can be distributed across distinct frequency bands, with band spacing ranging from GHz to THz, as presented in Fig. 4. To enable signal detection, the RARE needs to couple each band to a unique energy-level pair, denoted as |$\ket{3}\leftrightarrow\ket{n},n\in\{4,5,\cdots,N\}$. The electron transition between each energy-level pair down-converts the coupled band by its respective transition frequency $f_{\ket{3}\rightarrow\ket{n}}$, thereby simultaneously converting the entire multi-band RF signal to the baseband. Utilizing this principle, researchers have demonstrated the co-detection of 5 separate frequency bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [5].
Figure 4: Standard and multiband FDM-RAREs.
@@ IV-B Atomic-MIMO Communications
As a core technology in 5G and beyond, MIMO communications can greatly enhance the spectral efficiency through the diversity and multiplexing gains. As a representative, the recent study in [12] has demonstrated a good performance with an 20 Gbps 1500 m transmission distance in free-space-optical communications using advanced MIMO diversity combining codes. The advancement of MIMO thus motivates the research on atomic-MIMO communications. Its key idea is to employ an array of RAREs, in place of traditional antenna arrays, for signal detection. In this context, both the high sensitivity of Rydberg atoms and the high spectral efficiency of MIMO can be leveraged.
@@ IV-C RARE-empowered Quantum Wireless Sensing
In addition, RARE can also be utilized to sense angle-of-arrivals (AoA) and time-of-arrivals (ToA) for the purpose of localization. A self-heterodyne sensing technique was introduced in [14] to advance the current heterodyne sensing for high-resolution localization. This technique leverages the self-interference produced by the transmitted sensing signal as the reference signal, ensuring their waveforms to be identical. By doing so, it effectively overcomes the narrow instantaneous bandwidth limitation of RARE, enhancing the sensing bandwidth up to 100 MHz. Experiments show that a self-heterodyne RARE reduces the localization error by 20 dB compared with classic receivers.
@@ IV-D Quantum Many-body RAREs for External-noise Mitigation
One intriguing approach leverages the quantum many-body effect of Rydberg atoms to counterintuitively convert the power of external noise into that of wireless signal, achieving a noise-enhanced RARE [15]. In the approach, a collection of Rydberg atoms are interacted with each other. This many-body interaction transforms the original linear Schrödinger equation into a nonlinear one, which is shown to exhibit a bistability behavior. The incoming wireless signal could stimulate the system to switch between the two stable states. Then, by introducing a certain amount of external noise into the receiver, the switching magnitude can be amplified such that the output SNR is drastically improved. Experiment results demonstrate that the noise-enhanced RARE can surpass the sensitivity of traditional RAREs by 6.6 dB in the presence of external noise. This work opens a new promising direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the quantum many-body interactions.
V RARE for Multiband Sensing and Communication (RARE-MSAC)
@@
Figure 5: A single RARE is used to simultaneously detect communication signals at low and medium frequency bands, and sensing signals at high frequency bands.
@@
We leverage the advancement of RARE to introduce a novel application in PHY wireless systems——the RARE-MSAC. The key idea is to exploit the multi-level electron transitions of Rydberg atoms for detecting multiband signals with diverse sensing and communication (S&C) requirements. As illustrated in Fig. 5, a single RARE is employed. The atoms are prepared in an appropriate initial Rydberg state, |$\ket{3}$, where the transitions from |$\ket{3}$ to other Rydberg states, |$\ket{n},\forall n\geq 4$, can be coupled to distinct frequency bands. In this context, the RARE is capable of concurrently detecting communication signals at low ($<6{\rm GHz}$) or medium bands ($7{\rm GHz}-24{\rm GHz}$) to ensure coverage. Simultaneously, it can also capture sensing signals operating in the mmWave and THz bands to perform sensing tasks with varying granularity requirements, ranging from ToA estimation to heartbeat detection. It is amazing that all these S&C services are achieved by a single RARE device, which is nearly impossible for classic receivers due to the largely separated frequencies.
@@
Fig. 6 compares the MSAC performance of the RARE and classic receivers. We consider a dual-band S&C system. A sub-6G device transmits communication signals operating at $3.213\>{\rm GHz}$ with an instantaneous bandwidth of $500\>{\rm kHz}$. Simultaneously, a mmWave device transmits sensing signals operating at $30.618\>{\rm GHz}$ with a bandwidth of $100\>{\rm kHz}$. Similar to [4], the sensing signal is reflected by a target for detecting its vibration . A single RARE is deployed to detect these dual-band signals for providing S&C services. The Rydberg states are configured as |$\ket{3}=60D_{5/2}$, |$\ket{4}=61P_{3/2}$, and |$\ket{5}=62P_{3/2}$, respectively. The electron transition $60D_{5/2}\rightarrow 61P_{3/2}$ has a dipole moment of $2.04\times 10^{-26}\>{\rm C\cdot m}$ and resonates with the $3.213\>{\rm GHz}$ communication signal, while the transition $60D_{5/2}\rightarrow 62P_{3/2}$ has a dipole moment of $6.24\times 10^{-27}\>{\rm C\cdot m}$ and resonates with the $30.618\>{\rm GHz}$ sensing part. The power of the probe laser and coupling laser are set as $120{\rm\mu W}$ and $34\>{\rm mW}$, respectively. For comparison, two classic receivers are also deployed to independently complete the aforementioned S&C tasks, denoted as “CR1” for the sub-6G communication service and “CR2” for the mmWave sensing service in Fig. 6. The transmission power for the two signals increases from $10\>{\rm dBm}$ to $40\>{\rm dBm}$. As observed from the experimental results of Fig. 6, the spectrum efficiency in the communication service achieved by RARE is $\sim 6{\rm bps/Hz}$ higher than that of the classic receiver under Rayleigh fading channels. Additionally, the normalized mean square error (NMSE) of vibration sensing achieved by the RARE improves that achieved by the classic receiver by two orders of magnitude. Notably, these performance gains are accomplished by a single RARE without requiring additional RF circuits.
@@
Figure 6: Comparison of MSAC performance between a single RARE and two classic receivers, denoted as “CR 1” and “CR 2” respectively.

### v3

@@ Abstract
Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic wave measurement by harnessing the electron transition phenomenon of Rydberg atoms. Operating at the quantum scale, such receivers have the potential to breakthrough the performance limit of classic receivers, sparking a revolution in physical-layer wireless communications. The objective of this paper is to offer insights into RARE-empowered communication systems. We first provide a comprehensive introduction to the fundamental principles of RAREs. Then, a thorough comparison between RAREs and classic receivers is conducted in terms of the antenna size, sensitivity, and bandwidth. Subsequently, we overview the recent progresses in RARE-aided wireless communications, covering the frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques. Moreover, the unique application of RARE in multi-band communication and sensing is unveiled, followed by numerical experiments to show its performance superiority over classic receivers by $\sim 2.4\>{\rm bps/Hz}$ in spectrum efficiency and $\sim 7.2\>{\rm dB}$ in sensing accuracy. Finally, we conclude this paper by providing promising research directions.
@@ I Introduction
In current communication systems, receivers detect wireless signals by using an antenna to transform incident EM waves into electric currents [1]. This mechanism, however, presents insurmountable limitations to physical-layer (PHY) wireless communications. Firstly, the front-end circuits of a receiver, like amplifiers and mixers, introduce thermal noise into data currents, restricting the receiver’s sensitivity limit. In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its working bandwidth to a narrow range around the central frequency. To detect signals from sub-6G to THz, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band. Consequently, these challenges constrain the performance upper-bound of wireless communications [2].
@@
Recently, at the intersection of quantum sensing and wireless communication, Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3]. Rydberg atoms are highly-excited atoms with outer electrons far from the nucleus. Due to the large electric dipole moments, Rydberg atoms can strongly interact with incident EM waves, triggering electron transitions between energy levels. As a result, RAREs can accurately detect wireless signals by monitoring changes on the atoms’ quantum states resulting from electron transitions [2]. Representing a paradigm shift in receiver architecture to embrace quantum sensing, RAREs are gaining increasing attentions due to their quantum advantages. For instance, a RARE is immune to thermal noise as it involves no front-end circuits. Hence, its sensitivity limit can potentially exceed those of classic counterparts [4]. Moreover, a single RARE is able to interact with a large number of frequencies ranging from Megahertz (MHz) to THz band because of the abundant energy levels in Rydberg atoms [5]. These quantum advantages make RAREs a promising technology to revolutionize the next-generation wireless network.
@@
The objective of this paper is to provide a comprehensive guide for beginners to the RARE-empowered communication systems. To this end, we first introduce the fundamental characteristics of Rydberg atoms and the key components of a RARE. Next, the key indicators of RAREs are compared with classic receivers in terms of antenna size, sensitivity, and bandwidth. Subsequently, we review the state-of-the-art RARE techniques for wireless communications, covering frequency-division multiplexing (FDM), multiple-input-multiple-output (MIMO), quantum wireless sensing, and quantum many-body techniques. We then reveal a unique application of RAREs in multi-band communication and sensing systems—a single RARE can simultaneously detect communication signals at low frequencies to ensure coverage and sensing signals at high frequencies to enhance granularity. Numerical experiments confirm that RAREs outperform conventional counterparts, achieving a spectrum efficiency gain of $\sim 2.4\>{\rm bps/Hz}$ and a sensing accuracy improvement of $\sim 7.2\>{\rm dB}$. Last, we highlight promising future research directions on RARE from the PHY wireless communication perspective, including signal modeling, holographic atomic-MIMO, satellite communications, and related areas.
@@ II-A3 Properties of Rydberg atom
Given the high principal quantum number $n$, Rydberg atoms exhibit useful properties for sensing EM waves. First, the radius of an Rydberg atom is large due to its dependence on $n^{2}$ [2]. For instance, the atomic radius grows from $150a_{0}$ to $10^{4}a_{0}$ as $n$ increases from 101010 to 100100100, with $a_{0}=5.29\times 10^{-11}{\rm m}$ being the Bohr radius. The substantial increase in radius endows the Rydberg atom with a large electric dipole moment, thus making it very sensitive to external EM wave. Second, the transition frequency between adjacent Rydberg states is abundant as the energy level spacing scales as $n^{-3}$ [3]. For low energy levels ($n<20$), the transition frequency can reach the visible light frequency (>100>100> 100 THz). While for Rydberg states ($n>20$), the transition frequencies fall within the range of MHz to THz [4], exactly covering all the typical frequency bands used in PHY communications.
@@ II-B2 EIT based quantum measurement
Figure 2: EIT for measuring AM/FM/PM signals. For AM and PM detection, the detuning $\delta$ is zero, while for FM detection, $\delta$ is nonzero.
@@
  • The (1,2)(1,2)( 1 , 2 )-th element of the final quantum state, $\rho_{12}$, determines the rate at which Rydberg atoms absorb photons from the probe laser. The absorption loss of the probe-laser scales exponentially with $\rho_{12}$ [6]. Thereby, it is feasible to deduce the value of $\rho_{12}$ by measuring the output power of the probe laser using a photodetector (PD), as depicted in Fig. 1.
@@ II-B3 Receive signal processing
So far, we have discussed the mechanism of detecting AM/FM/PM wireless signals using RAREs. Recent studies have also reported the detection of multi-frequency signals [10], multi-user signals [8], and target sensing [4, 11], all based on the principles we have covered and will be elaborated in Section IV.
@@
The RARE technique represents a paradigm shift in wireless signal detection from using electronic circuits to embracing quantum sensing. This gives rise to a range of unique features for RARE-enabled PHY wireless communications, which are introduced in this section.
@@ III-A Antenna Size
According to the Chu limit, a lossless and impedance-matched antenna in a conventional receiver requires a physical dimension proportional to the wavelength, e.g. $\frac{\lambda}{2}$, so as to efficiently convert EM waves into signals on a transmission line. At low frequencies, the size of an antenna element can reach hundreds of centimeters, for example exceeding $150\>{\rm cm}$ when $f<100\>{\rm MHz}$, leading to bulky receiver architectures. Conversely, a RARE does not absorb energy from the field, but rather leverages the atom-light interaction to perform nondestructive sensing. This mechanism enables the development of electrically small Rydberg antennas that are significantly smaller than the wavelength. Typically, the geometry size of a Rydberg antenna—comprising an ensemble of Rydberg atoms—is on the order of $1\>{\rm cm}\times 1\>{\rm mm}$, with the vapor cell length around $1\>{\rm cm}$ and the laser diameter approximately $1\>{\rm mm}$. This wavelength-independent size offers substantial advantages, especially for low-frequency applications.
@@ III-C Bandwidth
In existing RARE studies, the term ”bandwidth” is often used to represent two distinct metrics. Firstly, it is utilized to define the range of frequencies that can trigger electron transitions of Rydberg atoms. This concept is commonly referred to as the “detectable frequency range” of a RARE in the literature. Secondly, it is also associated with the “instantaneous bandwidth”, which denotes the maximum data capacity that the receiver device can measure at any given moment. This term aligns with the standard definition of ”bandwidth” in communication engineering. Typically, a RARE is characterized by a broad detectable frequency range but a narrow instantaneous bandwidth.
@@ III-C2 Instantaneous bandwidth
Restricted by the long response time for an EIT process to become steady, the instantaneous bandwidth of a RARE is typically narrower than 10 MHz [7]. This bandwidth is comparable to that of 4G LTE systems (1$\sim$20 MHz). Recent efforts are endeavored to increase the RARE’s bandwidth. For example, researchers have proposed a spatiotemporal multiplexing technique of probe lasers to decrease the EIT response time (see [139] of [2]). This technique expands the RARE’s bandwidth to 100 MHz, approaching that of sub-6G systems specified by 5G standards (i.e., 20$\sim$100 MHz). To date, continuously improving RARE’s instantaneous bandwidth remains a hot topic.
@@ IV State-of-the-art RARE Techniques
There have been studies to develop advanced RARE techniques to enhance the communication and sensing performance. This section provides an overview of the most recent progresses of these state-of-the-art designs.
@@ IV-A Frequency-division Multiplexed RAREs
The FDM technology plays a crucial role in 5G wideband systems for mitigating multi-path interference. The advancement of FDM-RARE is regarded as the key enabler of RARE-empowered wideband communications. The state-of-the-art FDM-RAREs can be separated into two categories: standard and multi-band FDM-RAREs.
@@ IV-A1 Standard FDM-RARE
IV-A2 Multi-band FDM-RARE
The multi-band FDM-RARE leverages the broad detectable frequency range of a RAREs to serve as an integrated full-frequency communication platform. On this platform, signal frequencies can be distributed across distinct frequency bands, with band spacing ranging from GHz to THz, as presented in Fig. 4. To enable signal detection, the RARE needs to couple each band to a unique energy-level pair, denoted as |$\ket{3}\leftrightarrow\ket{n},n\in\{4,5,\cdots,N\}$. The electron transition between each energy-level pair down-converts the coupled band by its respective transition frequency $f_{\ket{3}\rightarrow\ket{n}}$, thereby simultaneously converting the entire multi-band RF signal to the baseband. Utilizing this principle, researchers have demonstrated the co-detection of 5 separate frequency bands (1.72, 12.11, 27.42, 65.11, and 115.75 GHz) with a 3-dB instantaneous bandwidth of 6 MHz, verifying RARE’s full-frequency sensing capability [5].
Figure 4: Standard and multi-band FDM-RAREs.
@@ IV-B Atomic-MIMO Communications
As a core technology in 5G and beyond, MIMO communications can greatly enhance the spectral efficiency through the diversity and multiplexing gains. The recent study in [13] has demonstrated a 20 Gbps 1500 m transmission in free-space-optical communications using MIMO diversity combining codes. The advancement of MIMO thus motivates the research on atomic-MIMO communications. Its key idea is to employ an array of RAREs, in place of traditional antenna arrays, for signal detection. In this context, both the high sensitivity of Rydberg atoms and the high spectral efficiency of MIMO can be leveraged.
@@ IV-C RARE-empowered Quantum Wireless Sensing
In addition, RARE can also be utilized to sense angle-of-arrivals (AoA) and time-of-arrivals (ToA) for the purpose of localization. A self-heterodyne sensing technique was introduced in [11] to advance the current heterodyne sensing for high-resolution localization. This technique leverages the self-interference produced by the transmitted sensing signal as the reference signal, ensuring their waveforms to be identical. By doing so, it effectively overcomes the narrow instantaneous bandwidth limitation of RARE, enhancing the sensing bandwidth up to 100 MHz. Experiments show that a self-heterodyne RARE reduces the localization error by orders of magnitude compared with classic receivers.
@@ IV-D Quantum Many-body RAREs for External-noise Mitigation
One intriguing approach leverages the quantum many-body effect of Rydberg atoms to counterintuitively convert the power of external noise into that of wireless signal, achieving a noise-enhanced RARE [15]. In the approach, a collection of Rydberg atoms are interacted with each other. This many-body interaction transforms the original linear Schrödinger equation into a nonlinear one, which is shown to exhibit a bistability behavior. The incoming wireless signal could stimulate the system to switch between the two stable states. Then, by introducing a certain amount of external noise into the receiver, the switching magnitude can be amplified such that the output SNR is drastically improved. Experiment results demonstrate that the noise-enhanced RARE can surpass the sensitivity of traditional RAREs by 6.6 dB in the presence of external noise. This work opens a new direction for advancing RARE: the detrimental impact of external noise can be turned to a beneficial effect through the quantum many-body interactions.
V RARE for Multi-band Sensing and Communication (RARE-MSAC)
@@
Figure 5: A single RARE is used to simultaneously detect communication and sensing signals at different frequency bands.
@@
We leverage the advancement of RARE to introduce a novel application in wireless systems——the RARE-MSAC. The key idea is to exploit the multi-level electron transitions of Rydberg atoms for detecting multi-band signals with diverse sensing and communication (S&C) requirements. As illustrated in Fig. 5, a single RARE is employed. The atoms are prepared in an appropriate initial Rydberg state, |$\ket{3}$, where the transitions from |$\ket{3}$ to other Rydberg states, |$\ket{n},\forall n\geq 4$, can be coupled to distinct frequency bands. In this context, the RARE is capable of concurrently detecting communication signals at low ($<6{\rm GHz}$) or medium bands ($7{\rm GHz}-24{\rm GHz}$) to ensure coverage. Simultaneously, it can also capture sensing signals operating in the mmWave and THz bands to perform sensing tasks with varying granularity requirements, ranging from ToA estimation to heartbeat detection. It is amazing that all these S&C services are achieved by a single RARE device, which is nearly impossible for classic receivers due to the largely separated frequencies.
@@
Fig. 6 compares the MSAC performance of the RARE and classic receivers. We consider a dual-band S&C system. A sub-6G device transmits communication signals operating at $3.213\>{\rm GHz}$ with an instantaneous bandwidth of $500\>{\rm kHz}$. Simultaneously, a mmWave device transmits sensing signals operating at $30.618\>{\rm GHz}$ with a bandwidth of $100\>{\rm kHz}$. Similar to [4], the sensing signal is reflected by a target for detecting its vibration. A single RARE is deployed to detect these dual-band signals. The Rydberg states are configured as |$\ket{3}=60D_{5/2}$, |$\ket{4}=61P_{3/2}$, and |$\ket{5}=62P_{3/2}$, respectively. The electron transition $60D_{5/2}\rightarrow 61P_{3/2}$ has a dipole moment of $2.04\times 10^{-26}\>{\rm C\cdot m}$ and resonates with the $3.213\>{\rm GHz}$ communication signal, while the transition $60D_{5/2}\rightarrow 62P_{3/2}$ has a dipole moment of $6.24\times 10^{-27}\>{\rm C\cdot m}$ and resonates with the $30.618\>{\rm GHz}$ sensing part. For comparison, two classic receivers are also deployed to independently complete the aforementioned S&C tasks, denoted as “CR1” for the communication service and “CR2” for the sensing. The transmission power increases from $-10\>{\rm dBm}$ to $10\>{\rm dBm}$. As observed in Fig. 6, the spectrum efficiency in the communication service achieved by RARE is $\sim 2.4{\rm bps/Hz}$ higher than that of CR1 under Rayleigh fading channels. Additionally, the normalized mean square error (NMSE) of vibration sensing achieved by the RARE improves that achieved by CR2 by 7.2$\sim$ 7.2\>∼ 7.2dB. Notably, these performance gains are accomplished by a single RARE without requiring additional RF circuits.
@@
Figure 6: Comparison of MSAC performance between a single RARE and two classic receivers.